OpenStudy (paystub-2):

-2= -(n-8) wat happens with the negative sign in front of the parantheses

OpenStudy (bahrom7893):

-2 = 8-n n = 8+2

OpenStudy (paystub-2):

wat happens with the negative sign in front of the parantheses?

OpenStudy (anonymous):

think of it as -1 does that help

OpenStudy (paystub-2):

but in the end does it all come out to be 10

OpenStudy (anonymous):

if there is a negative sign on both the sides, it gets reversed to positive in your above example -2 = -(n-8) can be written as 2 = n - 8 n = 10

OpenStudy (anonymous):

 does that help